### v1

@@ 4.1. Experiment Setup
Inc -v3 MI-FGSM 100* 45.2 42.3 36.7 13.5 12.9 6.3 7.9 9.8 30.5
NGI-MI-FGSM(ours) 100* 61.6 61.0 50.6 20.0 21.5 11.0 12.7 15.2 39.3
NI-FGSM 100* 49.1 47.7 39.0 14.0 13.0 6.1 8.0 11.5 32.0
NGI-NI-FGSM(ours) 100* 58.8 55.5 46.5 15.6 15.8 7.9 10.4 12.7 35.9
Inc -v4 MI-FGSM 55.0 99.7* 47.5 43.0 16.2 15.8 7.6 10.2 14.6 34.4
NGI-MI-FGSM(ours) 73.6 99.4* 63.5 54.1 26.3 25.8 16.5 16.5 22.8 44.3
NI-FGSM 60.6 100* 50.9 43.5 17.4 14.7 7.0 10.8 15.4 35.6
NGI-NI-FGSM(ours) 70.9 99.9* 61.4 51.4 18.0 17.4 9.6 13.1 17.9 40.0
IncRes -v2 MI-FGSM 58.4 50.3 97.7* 45.3 22.5 16.5 12.1 10.4 13.8 36.3
NGI-MI-FGSM(ours) 76.0 69.9 97.7* 64.2 36.0 28.1 24.1 15.7 20.7 48.0
NI-FGSM 63.1 54.7 99.0* 45.9 21.1 16.8 11.8 10.4 13.9 37.4
NGI-NI-FGSM(ours) 74.8 67.9 98.6* 57.7 25.8 19.8 14.3 12.3 14.8 42.9
Res -101 MI-FGSM 56.9 50.4 49.3 99.3* 24.8 21.7 13.0 11.3 13.8 37.8
NGI-MI-FGSM(ours) 71.5 66.4 64.4 99.3* 35.5 32.3 22.4 16.8 19.4 47.6
NI-FGSM 62.9 55.9 53.2 99.3* 25.7 22.2 12.3 12.4 15.4 39.9
NGI-NI-FGSM(ours) 70.2 64.6 64.3 99.2* 30.1 24.6 15.4 14.1 16.4 44.3
@@
4.2. Attack Evaluation
4.2.1.  Evaluation on Single-model Attacks
In this section, we initially investigate the vulnerability of neural networks in the single-model setting. Specifically, we choose a surrogate model to produce adversarial examples and assess their effectiveness across all seven selected models. We employ the attack success rate as our evaluation metric, which is the misclassification rate of corresponding models against adversarial examples. A higher value indicates that the adversarial examples exhibit greater transferability.
We compare with the chosen optimization-based and input transformation-based baselines. Moreover, we evaluate our method against recent state-of-the-art methods. The subsequent content elucidate the aforementioned experimental results in detail, respectively.
Attack on optimization-based attacks. Initially, we conduct attacks using optimization-based baselines, MI-FGSM and NI-FGSM. When combined with our approach, denoted as NGI-MI-FGSM and NGI-NI-FGSM, we assess their enhancement in attack performance over the baselines. Table 1 illustrates that our method boosts the attack success rate for all black-box models, with a significant improvement compared to MI-FGSM and NI-FGSM, respectively. In particular, NGI-MI-FGSM can typically enhance transferability by 11.1% to 19.3% under normally trained models. Conversely, for adversarially trained models the uplift lies between 4.7% and 13.5%. Moreover, the transferability of transformer-based models witnesses an average enhancement of approximately 6%, thereby further attesting to the efficacy and superiority of our approach.
@@ 4.2.1.  Evaluation on Single-model Attacks
The results reveal that our method effectively improves the attack success rate for both black-box and white-box models, particularly for black-box models, with an average improvement of 8%, 13.2%, and 10.7% compared to the DIM, TIM, and SIM baselines, respectively. Nevertheless, Lin et al. (Lin et al., 2019) discover that merging DIM, TIM, and SIM can further enhance the efficiency of optimization-based attacks. We regard this as one of the current state-of-the-art methods and will delve into it in Table 2.
Compare with state-of-the-art methods In this part, we compare our approach with several prominent advanced adversarial attacks that have been recently introduced. By integrating our proposed strategy with five top-performing methods, we examine their performance enhancement in black-box model attacks. As shown in Table 2, our approach can significantly enhance the transferability of the adversarial examples, outperforming the competing methods. For instance, when using Inc-Resv2 as the surrogate model for generating adversarial examples, our strategy achieves an average of 87.1%, 73.0%, 69.4%, 81.6%, and 80.5%, representing an average improvement of 3.0%, 4.7%, 5.1%, 5.5%, and 7.5% over SSA-SI-TI-DIM, SIA-MI-FGSM, BSR-MI-FGSM, Admix-SI-TI-DIM, and SI-TI-DIM, respectively. It’s worth noting that the recent RAP attack strategy, which claims to achieve superior black-box transferability with more steps, only reached an average attack success rate of 54.8% for RAP-SI-TI-DIM. Conversely, NGI-SI-TI-DIM attained a significantly higher average attack success rate of 76.2%, which substantiates the effectiveness of our approach.
@@ 4.2.2.  Evaluation on Ensemble-model Attacks
The strategy of employing ensemble models has been validated to significantly enhance the transferability of adversarial examples (Liu et al., 2017). In this section, we elucidate the effectiveness of our method within this attack paradigm. Specifically, we generate adversarial examples by attacking multiple surrogate models, e.g., Inc-v3, Inc-v4, Inc-Resv2, and Res-101, and subsequently evaluate their performance within chosen target models. Table 3 demonstrates that our strategy achieves an average attack success rate of 92.8%, 93.3%, 93.4%, 94.4%, and 94.3% under the ensemble-model setting. Further contrasting this with established methods like SI-TI-DIM, BSR-MI-FGSM, Admix-SI-TI-DIM, SSA-SI-TI-DIM, and SIA-MI-FGSM our approach exhibits a marked enhancement in transferability. Specifically, for adversarially trained models, the attack success rates range from 92.8% to 98.5%. Additionally, for models employing the transformer architecture, there is a remarkable increase in transferability, ranging from 4.0% to 12.1%.
@@
Attack HGD R&P NIPS-r3 JPEG FD Com-Defend RS Bit-Red NRP Avg.
SI-TI-DIM 90.4 86.3 86.7 91.5 89.1 90.0 68.6 76.0 75.1 83.7
NGI-SI-TI-DIM(ours) 94.7 93.8 93.7 96.3 93.8 94.9 78.5 84.2 86.9 90.8
Admix-SI-TI-DIM 90.8 87.1 87.3 92.0 88.3 89.4 68.1 73.7 73.8 83.4
NGI-Admix-SI-TI-DIM(ours) 95.3 93.9 93.5 96.5 92.7 93.9 78.5 84.9 86.5 90.6
SSA-SI-TI-DIM 95.9 95.4 95.3 96.8 95.1 96.2 90.3 89.5 90.7 93.9
NGI-SSA-SI-TI-DIM(ours) 97.5 97.1 97.4 98.4 97.0 97.5 93.2 91.9 92.2 95.8
@@ 4.2.3.  Evaluation on Defense Models
To thoroughly assess our approach, we evaluate its performance against nine recent adversarial defense strategies, including the top three from the NIPS defense competition (HGD (Liao et al., 2018), R&P (Xie et al., 2017), NIPS-r3), along with six advanced methods (Bit-Red (Xu et al., 2018), JPEG (Guo et al., 2017), FD (Liu et al., 2019), ComDefend (Jia et al., 2019), RS (Cohen et al., 2019), NRP (Naseer et al., 2020)). As depicted in Table 4, our approach exhibits strong attack performance against multiple defense strategies, notably the RS defense. While SI-TI-DIM and Admix-SI-TI-DIM achieve attack success rates of 68.6% and 68.1% against the RS defense, respectively, our strategy significantly enhances the transferability of adversarial examples, yielding improvements of 9.9% and 10.4% over these baselines. Furthermore, against the robust defense mechanism recently introduced as Neural Representation Purifier (NRP), our approach successfully attains an impressive attack success rate exceeding 86.5%. Moreover, our method consistently maintains remarkable attack success rates, averaging 90.8%, 90.6%, and 95.8%, further demonstrating its effectiveness.
@@ 4.3.1.  Effect of different modules
4.4. Effect of different input transformation strategy
In this section, we evaluate the effectiveness of different input transformation strategies within our multi-way attack strategy. To fully leverage the Neighborhood Gradient Information (NGI) without incurring additional time costs, our strategy involves increasing attack step in subsequent iterations to achieve sufficient perturbation. However, we observed that directly increasing the attack step often yields poor results on adversarially trained models, which typically have a more extensive attention region  (Dong et al., 2019). Additionally, our studies found that the MaskProcess strategy encourages models to focus more on non-discriminative regions. In the MM process, i.e., the multi-way attack phase, acquiring diverse gradient information effectively prevents overfitting that could result from directly increasing the step size. To validate whether MaskProcess focuses on non-discriminative regions, we conduct ablation experiments comparing the effects of different input transformation methods in multi-way attacks. Specifically, we compare the use of clean inputs, perturbed inputs (DI inputs), and MaskProcess inputs in one branch. The results are shown in Table 6. The results show that under normally trained models, the transferability of adversarial examples generated from MaskProcess inputs is comparable to that from DI inputs; however, under adversarially trained models, MaskProcess inputs significantly outperform DI inputs, showing an improvement of 2.2% to 3.8% in attack success rate. In conclusion, our multi-way attack strategy can significantly enhance the adversarial transferability across various models, particularly demonstrating advantages in adversarially trained models.
@@ 5.  Conclusions
Border Impact. In this study, we present a novel NGI-Attack, incorporating Example Backtracking and Multiplex Mask strategies. These methods exhibit substantial transferability potential under black-box attack scenarios, while concurrently emphasizing the inherent vulnerabilities of DNNs. This revelation highlights the necessity for intensive research aimed at enhancing network robustness, offering significant insights into building secure and dependable neural networks.

### v2

@@ 4.1. Experiment Setup
Inc -v3 MI-FGSM 100.0 45.2 42.3 36.7 13.5 12.9 6.3 7.9 9.8 30.5
NGI-MI-FGSM(ours) 100.0 61.6 61.0 50.6 20.0 21.5 11.0 12.7 15.2 39.3
NI-FGSM 100.0 49.1 47.7 39.0 14.0 13.0 6.1 8.0 11.5 32.0
NGI-NI-FGSM(ours) 100.0 58.8 55.5 46.5 15.6 15.8 7.9 10.4 12.7 35.9
VMI-FGSM 100.0 71.6 68.8 60.3 32.4 30.8 17.7 17.8 22.4 46.9
NGI-VMI-FGSM(ours) 100.0 82.4 79.0 73.2 46.0 43.6 29.5 23.9 29.4 56.3
Inc -v4 MI-FGSM 55.0 99.7 47.5 43.0 16.2 15.8 7.6 10.2 14.6 34.4
NGI-MI-FGSM(ours) 73.6 99.4 63.5 54.1 26.3 25.8 16.5 16.5 22.8 44.3
NI-FGSM 60.6 100.0 50.9 43.5 17.4 14.7 7.0 10.8 15.4 35.6
NGI-NI-FGSM(ours) 70.9 99.9 61.4 51.4 18.0 17.4 9.6 13.1 17.9 40.0
VMI-FGSM 77.4 99.7 70.6 62.5 38.7 36.9 23.6 22.9 31.0 51.5
NGI-VMI-FGSM(ours) 86.4 99.8 81.1 73.2 50.9 48.5 37.0 28.6 40.0 60.6
IncRes -v2 MI-FGSM 58.4 50.3 97.7 45.3 22.5 16.5 12.1 10.4 13.8 36.3
NGI-MI-FGSM(ours) 76.0 69.9 97.7 64.2 36.0 28.1 24.1 15.7 20.7 48.0
NI-FGSM 63.1 54.7 99.0 45.9 21.1 16.8 11.8 10.4 13.9 37.4
NGI-NI-FGSM(ours) 74.8 67.9 98.6 57.7 25.8 19.8 14.3 12.3 14.8 42.9
VMI-FGSM 79.0 72.9 97.8 68.3 47.8 40.4 33.7 23.8 28.2 54.7
NGI-VMI-FGSM(ours) 84.6 82.7 97.9 76.6 59.5 53.0 51.9 32.2 41.0 64.4
Res -101 MI-FGSM 56.9 50.4 49.3 99.3 24.8 21.7 13.0 11.3 13.8 37.8
NGI-MI-FGSM(ours) 71.5 66.4 64.4 99.3 35.5 32.3 22.4 16.8 19.4 47.6
NI-FGSM 62.9 55.9 53.2 99.3 25.7 22.2 12.3 12.4 15.4 39.9
NGI-NI-FGSM(ours) 70.2 64.6 64.3 99.2 30.1 24.6 15.4 14.1 16.4 44.3
VMI-FGSM 74.7 69.9 70.1 99.2 45.4 41.5 29.6 22.0 25.2 53.1
NGI-VMI-FGSM(ours) 83.4 79.2 79.8 99.4 59.0 53.7 44.6 27.0 32.2 62.0
Table 2: Attack success rates (%) of SGM on nine normally and adversarially trained models in the single-model setting. The adversarial examples are crafted on Res50 and DN121, respectively. * indicates the white-box model.
Res50 SGM 31.0 24.4 24.6 32.9 21.9 22.9 16.8 11.2 17.2 22.5
NGI-SGM 49.6 40.7 40.0 49.0 37.8 36.6 28.5 19.4 25.5 36.3
DN121 SGM 34.8 28.4 25.8 39.5 29.4 29.1 20.0 11.0 16.4 26.0
NGI-SGM 59.0 51.1 49.5 61.6 48.7 44.4 36.7 20.5 29.0 44.5
4.2. Attack Evaluation
4.2.1. Evaluation on Single-model Attacks
In this section, we initially investigate the vulnerability of neural networks in the single-model setting. Specifically, we choose a surrogate model to generate adversarial examples and assess their effectiveness across nine diverse target models, including normally trained and adversarially trained CNNs, as well as transformer-based architectures. We employ the Attack Success Rate (ASR) as our evaluation metric, which is the misclassification rate of corresponding models against adversarial examples. A higher value indicates that the adversarial examples exhibit greater transferability.
We compare with the chosen optimization-based and input transformation-based baselines. Moreover, we evaluate our method against recent state-of-the-art methods. The subsequent content elucidates the aforementioned experimental results in detail, respectively.
Attack on optimization-based attacks. Initially, we conduct attacks using optimization-based baselines, MI-FGSM, NI-FGSM, and VMI-FGSM. When combined with our approach, denoted as NGI-MI-FGSM, NGI-NI-FGSM, and NGI-VMI-FGSM, we assess their enhancement in attack performance over the baselines.
Table 1 illustrates that our method consistently boosts the ASR across all black-box models while maintaining comparable performance in white-box settings. Specifically, NGI increases the average ASR by 8.8%–11.7% when applied to MI-FGSM, 3.9%–5.5% for NI-FGSM, and up to 9.7% for VMI-FGSM, with no performance degradation in any scenario. More notably, under normally trained models, NGI-MI-FGSM achieves substantial improvements in transferability, ranging from 11.1% to 19.3%, while on adversarially trained models, the improvement remains consistent within 4.7%–13.5%. In addition, NGI also enhances cross-architecture transferability, with an average improvement of approximately 6.0% on transformer-based models such as ViT and Swin-B. These results collectively highlight the generalizability and superiority of our approach.
Attack on SGM. To further demonstrate the generality of NGI, we apply it to SGM, which improves adversarial transferability by suppressing gradients from residual blocks and instead emphasizes gradients from skip connections. We choose ResNet-50 and DenseNet-121 as surrogate models, as their architectures feature extensive skip connections, which are essential for the effectiveness of SGM.
As shown in Table 2, NGI-SGM achieves substantial improvements over the SGM baseline across all target models. Specifically, under the ResNet-50 surrogate, NGI-SGM raises the average ASR by 13.8%, while under DenseNet-121, the gain reaches 18.5% . These results further demonstrate the effectiveness of NGI as a general plug-in strategy across diverse attack paradigms.
@@ 4.2.1. Evaluation on Single-model Attacks
The results reveal that our method effectively improves the ASR for both black-box and white-box models, particularly for black-box models, with an average improvement of 8%, 13.2%, and 10.7% compared to the DIM, TIM, and SIM baselines, respectively. Nevertheless, Lin et al. (DBLP:journals/corr/abs-1908-06281) discover that merging DIM, TIM, and SIM can further enhance the efficiency of optimization-based attacks. We regard this as one of the current state-of-the-art methods and will delve into it in Table 3.
Compare with state-of-the-art methods In this part, we compare our approach with several prominent advanced adversarial attacks that have been recently introduced. By integrating our proposed strategy with five top-performing methods, we examine their performance enhancement in black-box model attacks. As shown in Table 3, our approach can significantly enhance the transferability of the adversarial examples, outperforming the competing methods. For instance, when using Inc-Resv2 as the surrogate model for generating adversarial examples, our strategy achieves an average of 87.1%, 73.0%, 69.4%, 81.6%, and 80.5%, representing an average improvement of 3.0%, 4.7%, 5.1%, 5.5%, and 7.5% over SSA-SI-TI-DIM, SIA-MI-FGSM, BSR-MI-FGSM, Admix-SI-TI-DIM, and SI-TI-DIM, respectively. It’s worth noting that the recent RAP attack strategy, which claims to achieve superior black-box transferability with more steps, only reached an average ASR of 54.8% for RAP-SI-TI-DIM. Conversely, NGI-SI-TI-DIM attained a significantly higher average ASR of 76.2%, which substantiates the effectiveness of our approach.
@@ 4.2.2. Evaluation on Ensemble-model Attacks
The strategy of employing ensemble models has been validated to significantly enhance the transferability of adversarial examples (DBLP:conf/iclr/LiuCLS17). In this section, we elucidate the effectiveness of our method within this attack paradigm. Specifically, we generate adversarial examples by attacking multiple surrogate models, e.g., Inc-v3, Inc-v4, Inc-Resv2, and Res-101, and subsequently evaluate their performance within chosen target models. Table 4 demonstrates that our strategy achieves an average ASR of 92.8%, 93.3%, 93.4%, 94.4%, and 94.3% under the ensemble-model setting. Further contrasting this with established methods like SI-TI-DIM, BSR-MI-FGSM, Admix-SI-TI-DIM, SSA-SI-TI-DIM, and SIA-MI-FGSM our approach exhibits a marked enhancement in transferability. Specifically, for adversarially trained models, the ASR ranges from 92.8% to 98.5%. Additionally, for models employing the transformer architecture, there is a remarkable increase in transferability, ranging from 4.0% to 12.1%.
@@
Attack HGD R&P NIPS-r3 JPEG FD Com-Defend RS Bit-Red NRP DiffPure Avg.
SI-TI-DIM 90.4 86.3 86.7 91.5 89.1 90.0 68.6 76.0 75.1 61.6 81.5
NGI-SI-TI-DIM(ours) 94.7 93.8 93.7 96.3 93.8 94.9 78.5 84.2 86.9 69.8 88.7
Admix-SI-TI-DIM 90.8 87.1 87.3 92.0 88.3 89.4 68.1 73.7 73.8 61.0 81.2
NGI-Admix-SI-TI-DIM(ours) 95.3 93.9 93.5 96.5 92.7 93.9 78.5 84.9 86.5 71.7 88.7
SSA-SI-TI-DIM 95.9 95.4 95.3 96.8 95.1 96.2 90.3 89.5 90.7 84.3 93.0
NGI-SSA-SI-TI-DIM(ours) 97.5 97.1 97.4 98.4 97.0 97.5 93.2 91.9 92.2 89.6 95.2
@@ 4.2.3. Evaluation on Defense Models
To thoroughly assess our approach, we evaluate its performance against ten recent adversarial defense strategies, including the top three from the NIPS defense competition (HGD (DBLP:conf/cvpr/LiaoLDPH018), R&P (xie2017mitigating), NIPS-r3333https://github.com/anlthms/nips-2017/tree/master/mmd), along with seven advanced methods (Bit-Red (DBLP:conf/ndss/Xu0Q18), JPEG (guo2017countering), FD (liu2019feature), ComDefend (DBLP:conf/cvpr/JiaWCF19), RS (DBLP:conf/icml/CohenRK19), NRP (DBLP:conf/cvpr/NaseerKHKP20)), DiffPure (nie2022DiffPure). As depicted in Table 5, our approach exhibits strong attack performance against multiple defense strategies, notably the RS defense. While SI-TI-DIM and Admix-SI-TI-DIM achieve ASR of 68.6% and 68.1% against the RS defense, respectively, our strategy significantly enhances the transferability of adversarial examples, yielding improvements of 9.9% and 10.4% over these baselines. Furthermore, against the robust defense mechanism recently introduced as Neural Representation Purifier (NRP), our approach successfully attains an impressive ASR exceeding 86.5%. We further assess our approach under the recent state-of-the-art adversarial purification method, DiffPure. Despite DiffPure’s robust defense, our strategy achieves notable attack success rates of 69.8%, 71.7%, and 89.6%. Moreover, our method consistently maintains remarkable attack success rates, averaging 88.7%, 88.7%, and 95.2%, further demonstrating its effectiveness.
@@ 4.3.1. Effect of different modules
4.3.2. Effect of different input transformation strategies
In this section, we evaluate the effectiveness of different input transformation strategies within our multi-way attack strategy. To fully leverage the Neighborhood Gradient Information (NGI) without incurring additional time costs, our strategy involves increasing attack step in subsequent iterations to achieve sufficient perturbation. However, we observed that directly increasing the attack step often yields poor results on adversarially trained models, which typically have a more extensive attention region  (dong2019evading). Additionally, our studies found that the MaskProcess strategy encourages models to focus more on non-discriminative regions. In the MM process, i.e., the multi-way attack phase, acquiring diverse gradient information effectively prevents overfitting that could result from directly increasing the step size. To validate whether MaskProcess focuses on non-discriminative regions, we conduct ablation experiments comparing the effects of different input transformation methods in multi-way attacks. Specifically, we compare the use of clean inputs, perturbed inputs (DI inputs), and MaskProcess inputs in one branch. The results are shown in Table 7. The results show that under normally trained models, the transferability of adversarial examples generated from MaskProcess inputs is comparable to that from DI inputs; however, under adversarially trained models, MaskProcess inputs significantly outperform DI inputs, showing an improvement of 2.2% to 3.8% in attack success rate. In conclusion, our multi-way attack strategy can significantly enhance the adversarial transferability across various models, particularly demonstrating advantages in adversarially trained models.
@@ 4.3.4. Multiplex Mask Probability $P$
Table 8: Comparison of average ASR with standard deviation (%) on nine target models. Each result is averaged over five random trials using Inception-v3 as the surrogate model.
[…]
MI-FGSM 100.00 ($\pm$ 0.00) 44.34 ($\pm$ 0.46) 41.56 ($\pm$ 0.42) 36.16 ($\pm$ 0.33) 14.00 ($\pm$ 0.33) 12.82 ($\pm$ 0.58) 6.22 ($\pm$ 0.31) 7.88 ($\pm$ 0.34) 10.12 ($\pm$ 0.53)
NGI-MI-FGSM (ours) 100.00 ($\pm$ 0.00) 62.82 ($\pm$ 0.75) 59.34 ($\pm$ 0.41) 50.16 ($\pm$ 0.93) 21.06 ($\pm$ 0.48) 20.62 ($\pm$ 0.60) 11.16 ($\pm$ 0.42) 11.54 ($\pm$ 0.51) 15.58 ($\pm$ 0.31)
Table 9: Bootstrap hypothesis test (10,000 resamples, 5 independent runs) comparing the ASR of NGI-MI-FGSM with its baseline MI-FGSM on eight black-box target models. One-sided $p$-values are reported; a $p$-value below 0.05 indicates that the improvement of NGI-MI-FGSM is statistically significant.
Model Bootstrap p-value Significant (p <0.05)
Inc-v4 0.0057 TRUE
Inc-Res-v2 0.0038 TRUE
Res-101 0.005 TRUE
Inc-v3ens3 0.001 TRUE
Inc-v3ens4 0.002 TRUE
IncRes-v2ens 0.0018 TRUE
Vit 0.0001 TRUE
Swin-b 0.0015 TRUE
4.4. Statistical Robustness and Significance Validation
To enhance the credibility and statistical robustness of our evaluation, we perform five independent runs for each experimental setting using different random seeds. As shown in Table 8, we report the average ASR along with the standard deviation on eight black-box models when using Inception-v3 as the surrogate model.
We observe that attacks incorporating NGI-Attack exhibit slightly higher standard deviations compared to the baseline methods. This is primarily due to the stochastic nature of the Multiplex Mask strategy adopted in the second stage of our NGI-Attack. The random masking of inputs introduces greater diversity in gradient information, which improves transferability but also increases variance across trials. Nonetheless, NGI consistently outperforms the baseline in terms of average ASR, validating its effectiveness.
To further validate the statistical significance of our improvements, we conduct a bootstrap hypothesis test (bootstrap) with 10,000 resampling iterations. Specifically, we compare the ASR results of Table 8 and calculate one-sided p-values for each black-box target model. As presented in Table 9, all p-values are below the 0.05 threshold, confirming that the improvement of NGI-MI-FGSM over MI-FGSM is statistically significant. This analysis provides strong empirical evidence supporting the robustness and reliability of our method.
@@ 5. Conclusions
Limitation. While NGI-Attack significantly enhances adversarial transferability, it still exhibits certain limitations. First, its performance may degrade when there is a substantial architectural gap between the surrogate and target models, as the NGI may fail to align well with the target model’s decision boundaries. Second, a significant domain gap between source and target data can also reduce the effectiveness of the attack. These limitations suggest promising directions for improving the robustness and generalizability of NGI-based attacks.
Broader Impact. In this study, we present a novel NGI-Attack, incorporating Example Backtracking and Multiplex Mask strategies. These methods exhibit substantial transferability potential under black-box attack scenarios, while concurrently emphasizing the inherent vulnerabilities of DNNs. This revelation highlights the necessity for intensive research aimed at enhancing network robustness, offering significant insights into building secure and dependable neural networks.
Future Work. Moreover, since NGI-Attack is fundamentally a gradient-based strategy, its core principles are not restricted to vision models. Future work could explore its extension to other modalities such as natural language processing and speech, where gradients also play a central role in model behavior.